Vertical fluctuations were measured at the ATF-DR by the laser wire in present and last weeks. The vertical positions were calculated by laser-compton intensies per 1mA of beam current at off peaks, since the beams were in accumulating mode, i.e. the beam intensity decreased during the measurements. Typical vertical beam size was 6um at the laser wire. Resultant fluctuations were 2um and 1um of peak-to-peak with a period of 200~400sec which were observed at 7am and 7pm, respectively. In order to see any intensity fluctuations, the laser-compton intensities at the peak were measured. No apparent fluctuation was observed. Before this summer shutdown (two years ago), the fluctuation was 5um. So, we may see improvements of the cooling system. Also, a 100Hz fluctuation was confirmed to be still there. To compensate the 100Hz fluctuation, a ripple filter was installed before this beam shift. However, a fuse of the filter was blew out. So, we could not check the filter effect.
Following suggestions were discussed. The measurements shall be in the re-filling mode in order to reduce possible beam-intensity dependance. Temperature shall be monitored at the same time for its correlation. In December, many temperature monitors will be installed at the DR for safety operation of multi-bunch beams.
Major updates are (1) CFRP of the reference bar instead of Aluminum for lower thermal expansion, (2) thicker granite table ( from 20cm to 50cm), 100 x 150 x 50cm3, 2.2 tons, removing iron-frame for stability.
For the CFRP reference bar, concerns of hygroscopic property and anisotropy of the thermal expansion were pointed out. Also, it was suggested that four legs supporting the reference bar should be replaced with two plates for more rigid support where a hole is necessary for beam. In present design, the granite table is put on iron channels with ball bearing movers, and the channels are set on four level jacks. Since we can put the granite table horizontally in good position, the channels with the movers may be removed. Optical anchor system requires slightly extended reference bar and holes in the granite table for laser beams.
The system was analyzed by FEM model in the same way as before. While the deformation due to self-weight was estimated to be 81um (20um) on top of the reference frame, the relative one was 1.4um (4.5um) between the center and both ends, where values in parentheses@are those in the previous design. Resonance frequencies were also calculated to be 27, 40, 42 and 54Hz at 1st, 2nd, 3rd and 4th, which are lower than the previous design. Since the decrease of the frequencies is due to weaker support with the channels, the above improvement must recover them. Good support system would have the 1st resonance frequency above 100Hz.
The LGS files of the LLNL system have been successfully read and they produced CAD figures at KEK. Since these data are huge for our computer power, approximate structures will be used for ANSYS analysis.
J. Frisch has presented the optical anchor system for the nanoBPM at the 2nd mini-workshop. The LLNL system should be supported by three point springs with active feedbacks of electro-static actuator, over whose tops seismometers are set. Each three point has a mirror for laser interferometer with two vacuum paths which may cancel horizontal displacement, since the paths are not vertical.
The original concept has been proposed for a support system of final quadrupole (Q) magnets in a LC detector by M. Woods at SLAC. Basic assumption is an existence of a bedrock for a fixed pivot of laser interferometer. The two Q magnets shall be stabilized against the bedrock by laser interferometers. Prototype optics of optical anchor has also been proposed by S.Myers at SLAC. With a L shaped beam splitter put on a Q magnet, vertical displacements can be stabilized independently from horizontal ones if an injection angle of laser beam is 60 degrees . Although the prototype optics can get an interference between two vertical positions, it must be difficult to realize it in the LC detector.
The concept of optical anchor has been verified with a 10m arm interferometer at SLAC, where the stability of 0.2nm rms has been demonstrated. After this success, the optical anchor R&D has been moved to University of British Columbia (UBC). The UBC team constructed the prototype system which can control displacement with a heavy load. They also succeeded to get sub-nm stability with a close feedback loop. However, none of them verified stabilization of vertical displacements.
For the nanoBPM system, one of laser arm would be set on floor and the other arm shall be "vertically" set. Therefore, a key issue must be isolation/cancellation of vertical displacement from horizontal one. Higashi will design an optics for this purpose.